# Sharp Eyes and Memory for VideoLLMs: Information-Aware Visual Token Pruning for Efficient and Reliable VideoLLM Reasoning

### v1

@@ Methods
Video Large Language Models (VideoLLMs) have demonstrated remarkable capabilities in video understanding and reasoning tasks (Li et al. 2023, 2024b; Zhang et al. 2024b; Lin et al. 2023). However, the extensive spatiotemporal information in long videos consumes a significant portion of LLM input tokens (Li et al. 2024a; Xu et al. 2024a), leading to excessive computational overhead and memory usage (Xu et al. 2024b). As shown in Table 1, existing pre-LLM stage approaches have explored visual token pruning at various granularity levels to reduce token count and memory consumption. However, these methods (Chen et al. 2024; Lin et al. 2025; Hyun et al. 2025; Tao et al. 2025; Fu et al. 2024) uniformly apply fixed pruning ratios, which typically represent local optima tuned for specific datasets, lacking adaptive pruning based on video information content. This limitation affects their generalization and robustness. Moreover, since the primary objective of visual pruning is computational efficiency, the pruning process itself should maintain minimal complexity to avoid additional overhead. Existing approaches that rely on computationally intensive clustering algorithms (Huang, Zhou, and Han 2024) or mathematically sophisticated planning techniques (Sun et al. 2025; Alvar et al. 2025) may inadvertently introduce non-trivial overhead to the pruning process shown in Table 1, potentially limiting their practical applicability in real-world deployment scenarios where computational efficiency is paramount. To address these challenges, in pre-LLM stage, we propose a novel adaptive pruning method (Visual SharpV) that computes spatiotemporal importance for each token and determines frame-specific pruning ratios based on information volume, achieving fine-grained information preservation while maintaining low complexity.
@@
In the intra-LLM stage, as demonstrated in Table 1, most existing works reduce cache by analyzing exposed attention scores. For instance, (Chen et al. 2024; Lin et al. 2025) retain visual tokens with high attention scores in specific layers based on inefficient visual attention phenomena, while (Huang, Zhou, and Han 2024) preserves tokens with high cross-attention scores. Although attention-based analysis provides reasonable motivation and valuable insights into attention mechanism dynamics, the emergence of efficient attention computation methods like flash attention (Dao et al. 2022; Dao 2023), which reduces the quadratic complexity of native attention to linear complexity while not exposing attention scores, renders such approaches incompatible. Therefore, we propose an information-theoretic perspective for intra-LLM stage pruning by measuring similarity between decoder layer inputs and original visual features to dynamically retain visual tokens for subsequent generation steps. Our method reduces complexity from $O(n^{2}\cdot d)$ to $O(n\cdot d)$, fulfilling the lightweight design principle and the core objective of pruning.
@@
We integrate SharpV into two widely adopted VideoLLMs with varying capabilities: PLLaVA (Xu et al. 2024a), LLaVA-OneVision (Li et al. 2024a). Extensive evaluations are conducted on four public datasets covering various video types (long, medium, short), including MVBench (Li et al. 2024c), VideoMME (Fu et al. 2025), NExTQA(Xiao et al. 2021), ActNet-QA(Yu et al. 2019), with comparisons against recent SOTA methods and baselines. Our experiments demonstrate the effectiveness of information-aware adaptive pruning ratios: at similar or lower retention ratios, SharpV achieves higher accuracy than SOTA methods while delivering more significant speedups due to its extremely low computational overhead.
@@
Videos inherently exhibit spatio-temporal locality, which has been exploited in deep learning approaches. Early works like 3D CNNs (Carreira and Zisserman 2017; Tran et al. 2015) and VideoTransformers (Tong et al. 2022; Feichtenhofer et al. 2022) achieved promising results in feature learning by leveraging this redundancy. Similarly, in multimodal LLM inference scenarios, visual token compression methods have enabled efficient model reasoning. Some works employ various strategies to identify spatially important tokens, typically by filtering tokens based on features from visual encoders (Rao et al. 2021; Vasu et al. 2023; Bolya et al. 2022; Kong et al. 2022). For instance, PruneMerge (Shang et al. 2024) selects crucial tokens based on inherent attention characteristics (Radford et al. 2021) between [CLS] token and other visual token. Other works rely on similarity-based methods to identify important tokens. DivPrune (Alvar et al. 2025) reduces spatial redundancy by solving for the maximally diverse set that maintains minimum pairwise distances. Cross-modal interaction methods like SimIgnore (Zhang et al. 2025) filter instruction-relevant visual tokens by comparing their similarity with prompt tokens. However, these approaches only consider spatial redundancy while neglecting temporal redundancy in videos.
For spatio-temporal video compression, LLaVA-Scissor (Sun et al. 2025) preserves semantic information through two-stage connectivity analysis, but suffers from high computational costs due to graph construction and Union-Find algorithms. PruneVid (Huang, Zhou, and Han 2024) focuses on compressing static backgrounds but requires video segmentation via clustering, which is computationally intensive and struggles with dynamic parameter settings. VidCom2 (Liu et al. 2025) quantifies frame uniqueness through global token representations, which produces excessively coarse-grained features and proves inadequate for real-time inference scenarios. Dynamic methods like DyCoKe (Tao et al. 2025) and Frame Fusion (Fu et al. 2024) efficiently compress inter-frame tokens through similarity measures, yet face challenges in setting adaptive thresholds for diverse video types.
@@
During prefilling and decoding phases of multimodal large models, most approaches reduce computational costs and KV cache by analyzing attention distributions. FastV (Chen et al. 2024) and VTW (Lin et al. 2025) discovered that visual tokens receive significantly diminished attention in deeper layers, enabling pruning or removal in late stages. FitPrune (Ye et al. 2025) formulates pruning as a statistical distribution fitting problem, minimizing attention distribution discrepancies pre- and post-pruning. PrunerVid (Huang, Zhou, and Han 2024) performs uniform layer-wise pruning based on cross-attention scores between prompts and visual tokens, effectively reducing KV cache. DyCoKe (Tao et al. 2025) designs a special DP Cache to retain potentially useful tokens according to attention scores during subsequent decoding stages, optimizing matrix operations. Frame Fusion (Fu et al. 2024) accumulates cross-layer attention scores to measure token importance.
While attention-based pruning is theoretically reasonable, modern hardware acceleration strategies like flash attention (Dao et al. 2022; Dao 2023) have reduced quadratic complexity to linear complexity at the cost of hiding attention scores, rendering many existing pruning methods obsolete. Addressing this challenge, SharpV pioneers a novel pruning paradigm that operates independently of exposed attention scores, providing fresh insights for visual token compression.
@@
As is well-known, cosine similarity effectively measures directional alignment between vectors. In this work, we aim to quantify token dissimilarity as a positive measure. While some approaches simply rescale cosine similarity using (1 - cosine similarity), this becomes problematic in high-dimensional spaces (Tessari, Yao, and Hogan 2024; Yamagiwa, Oyama, and Shimodaira 2024), where random vectors have expected cosine similarity approaching zero (Steck, Ekanadham, and Kallus 2024).
@@
While some approaches compute pairwise token similarity to preserve critical semantic information (Alvar et al. 2025; Sun et al. 2025), this inevitably leads to quadratic computational complexity. Unlike existing methods (Liu et al. 2025) that coarsely estimate frame-level uniqueness, we propose a novel token-level spatio-temporal importance assessment mechanism. Assisted by our designed Dissimilarity Computation Module, each token’s importance is comprehensively evaluated along dual dimensions.
@@
As shown in Fig. 3, we comprehensively compare the cosine similarity between hidden states $G_{l}=(S_{l},V_{l},I_{l})\in\mathbb{R}^{(N_{s}+N_{v}+N_{p})\times d}$ at different layers $l$ and their original positions in $G=(S,V,I)\in\mathbb{R}^{(N_{s}+N_{v}+N_{p})\times d}$ for both LLaVA-OneVision-7B and PLLaVA, visualizing their average values. A striking observation is that in both models, the visual tokens $V_{l}$ exhibit high similarity on average in shallow layers and demonstrate a clear decreasing trend to their original counterparts $V$ as layer depth increases. In contrast, system and instruction tokens rapidly converge to near-zero similarity (0-0.2 range). Mathematically, the similarity of random high-dimensional vectors approaches zero, suggesting that system and instruction tokens undergo more complex unknown transformations through Attention and MLP operations. Meanwhile, visual tokens show measurable degradation of original visual information via cosine similarity, which we term Visual Information Degradation. Although the exact relationships between transformed token embeddings remain unclear (Zhang et al. 2025), the similarity patterns reveal that visual information undergoes rapid decay in shallow layers resembling human memory curves before stabilizing at low similarity levels (¡0.2) after layer 14. This suggests LLMs primarily process visual information in early layers, with diminishing importance in deeper layers.
@@
Research on visual attention reveals a consistent pattern: Inefficient Visual Attention phenomenon (Chen et al. 2024; Lin et al. 2025) validate that visual tokens receive drastically diminished attention scores in deeper layers of multimodal LLMs. We further quantify this via cosine similarity, a robust metric for high-dimensional feature decay (Kalhori, Alamdari, and Ye 2018; Luo et al. 2018; Dehak et al. 2010; Lahitani, Permanasari, and Setiawan 2016).
From the perspective of information theory, this degradation pattern aligns with the concept of progressive information compression in hierarchical systems. The rapid similarity decay in early layers suggests intensive feature extraction, while the stable low similarity in deeper layers indicates visual information has been sufficiently encoded into subsequent representations. This mirrors the information bottleneck principle (Tishby, Pereira, and Bialek 2000; Hsu, Kennedy, and Chang 2006), where networks discard irrelevant details while preserving task-relevant features through successive transformations (Tishby and Zaslavsky 2015). The distinct behavior of visual versus textual tokens further implies modality-specific information (Yang et al. 2022; Barsalou et al. 2003) processing pathways in multimodal LLMs, analogous to attentional load theory in human perception (Chun and Wolfe 2005).
@@
We employ lmms-eval (Zhang et al. 2024a) as our primary evaluation framework. For PLLaVA (Xu et al. 2024a) specifically, we use its official implementation to ensure fair comparison. Our comprehensive evaluation covers diverse video benchmarks, including MVBench (Li et al. 2024c) and VideoMME (Fu et al. 2025) for multi-dimensional video understanding and NExT-QA (Xiao et al. 2021) and ActNet-QA (Yu et al. 2019) for video question answering.
@@
We implement FastV (Chen et al. 2024), PruMerge (Shang et al. 2024) and DyCoKe (Tao et al. 2025) on two widely used LLMs for video understanding: PLLaVA(Xu et al. 2024a) and LLaVA-OneVision (Li et al. 2024a) with details shown in Tab.2. To demonstrate scalability, we evaluate 0.5B and 7B versions of LLaVA-OneVision.
@@
The term Token Budget combines the visual pruning ratio (VR) from the pre-LLM stage and the memory pruning ratio (MR) from the intra-LLM stage, defined as $VR\times MR$ following (Tao et al. 2025; Fu et al. 2024). VR measures the ratio of retained visual tokens fed into LLMs (Liu et al. 2025; Alvar et al. 2025), while MR estimates KV-Cache utilization during decoding (Fu et al. 2024). For efficiency comparison, we evaluate FLOPs, GPU memory and acceleration metrics: Time To First Token (TTFT) and Time Per Output Token (TPOT).
@@ Conclusion
SharpV is the first framework to explore intra-LLM visual token pruning without relying on exposed attention scores, achieving linear complexity and compatible with Flash Attention. Furthermore, we propose a simple yet effective pre-LLM token selection method, offering a self-calibrated, adaptive thresholding scheme with strong generality and robustness based on information measurement. These components synergize to reduce the final pruning rate around 12%, while frequently improving accuracy across multiple benchmarks due to reduced visual noise, enhancing both efficiency and reliability in Video-LLM reasoning.

### v2

@@ Methods
Video Large Language Models (VideoLLMs) have demonstrated remarkable capabilities in video understanding and reasoning tasks (videochat; llavanextvideo; llavavideo; videollava). However, the extensive spatiotemporal information in long videos consumes a significant portion of LLM input tokens (llavaov; pllava), leading to excessive computational overhead and memory usage (slowfast). As shown in Table 1, existing pre-LLM stage approaches have explored visual token pruning at various granularity levels to reduce token count and memory consumption. However, these methods (fastv; vtw; sttm; dycoke; framefusion) uniformly apply fixed pruning ratios, which typically represent local optima tuned for specific datasets, lacking adaptive pruning based on video information content. This limitation affects their generalization and robustness. Moreover, since the primary objective of visual pruning is computational efficiency, the pruning process itself should maintain minimal complexity to avoid additional overhead. Existing approaches that rely on computationally intensive clustering algorithms (prunevid) or mathematically sophisticated planning techniques (llavascissor; divprune) may inadvertently introduce non-trivial overhead to the pruning process shown in Table 1, potentially limiting their practical applicability in real-world deployment scenarios where computational efficiency is paramount. To address these challenges, in pre-LLM stage, we propose a novel adaptive pruning method (Visual SharpV) that computes spatiotemporal importance for each token and determines frame-specific pruning ratios based on information volume, achieving fine-grained information preservation while maintaining low complexity.
@@
In the intra-LLM stage, as demonstrated in Table 1, most existing works reduce cache by analyzing exposed attention scores. For instance, (fastv; vtw) retain visual tokens with high attention scores in specific layers based on inefficient visual attention phenomena, while (prunevid) preserves tokens with high cross-attention scores. Although attention-based analysis provides reasonable motivation and valuable insights into attention mechanism dynamics, the emergence of efficient attention computation methods like flash attention (flashattention; flashattention2), which reduces the quadratic complexity of native attention to linear complexity while not exposing attention scores, renders such approaches incompatible. Therefore, we propose an information-theoretic perspective for intra-LLM stage pruning by measuring similarity between decoder layer inputs and original visual features to dynamically retain visual tokens for subsequent generation steps. Our method reduces complexity from $O(n^{2}\cdot d)$ to $O(n\cdot d)$, fulfilling the lightweight design principle and the core objective of pruning.
@@
We integrate SharpV into two widely adopted VideoLLMs with varying capabilities: PLLaVA (pllava), LLaVA-OneVision (llavaov). Extensive evaluations are conducted on four public datasets covering various video types (long, medium, short), including MVBench (mvbench), VideoMME (videomme), NExTQA(nextqa), ActNet-QA(activitynet), with comparisons against recent SOTA methods and baselines. Our experiments demonstrate the effectiveness of information-aware adaptive pruning ratios: at similar or lower retention ratios, SharpV achieves higher accuracy than SOTA methods while delivering more significant speedups due to its extremely low computational overhead.
@@
Videos inherently exhibit spatio-temporal locality, which has been exploited in deep learning approaches. Early works like 3D CNNs (3dcnn1; 3dcnn2) and VideoTransformers (tong2022videomae; spatiotemporallearners) achieved promising results in feature learning by leveraging this redundancy. Similarly, in multimodal LLM inference scenarios, visual token compression methods have enabled efficient model reasoning. Some works employ various strategies to identify spatially important tokens, typically by filtering tokens based on features from visual encoders (dynamicvit; fastvit; tome; spvit). For instance, PruneMerge (prumerge) selects crucial tokens based on inherent attention characteristics (clipattention) between [CLS] token and other visual token. Other works rely on similarity-based methods to identify important tokens. DivPrune (divprune) reduces spatial redundancy by solving for the maximally diverse set that maintains minimum pairwise distances. Cross-modal interaction methods like SimIgnore (simignore) filter instruction-relevant visual tokens by comparing their similarity with prompt tokens. However, these approaches only consider spatial redundancy while neglecting temporal redundancy in videos.
For spatio-temporal video compression, LLaVA-Scissor (llavascissor) preserves semantic information through two-stage connectivity analysis, but suffers from high computational costs due to graph construction and Union-Find algorithms. PruneVid (prunevid) focuses on compressing static backgrounds but requires video segmentation via clustering, which is computationally intensive and struggles with dynamic parameter settings. VidCom2 (vicom2) quantifies frame uniqueness through global token representations, which produces excessively coarse-grained features and proves inadequate for real-time inference scenarios. Dynamic methods like DyCoKe (dycoke) and Frame Fusion (framefusion) efficiently compress inter-frame tokens through similarity measures, yet face challenges in setting adaptive thresholds for diverse video types.
@@
During prefilling and decoding phases of multimodal large models, most approaches reduce computational costs and KV cache by analyzing attention distributions. FastV (fastv) and VTW (vtw) discovered that visual tokens receive significantly diminished attention in deeper layers, enabling pruning or removal in late stages. FitPrune (fitprune) formulates pruning as a statistical distribution fitting problem, minimizing attention distribution discrepancies pre- and post-pruning. PrunerVid (prunevid) performs uniform layer-wise pruning based on cross-attention scores between prompts and visual tokens, effectively reducing KV cache. DyCoKe (dycoke) designs a special DP Cache to retain potentially useful tokens according to attention scores during subsequent decoding stages, optimizing matrix operations. Frame Fusion (framefusion) accumulates cross-layer attention scores to measure token importance.
While attention-based pruning is theoretically reasonable, modern hardware acceleration strategies like flash attention (flashattention; flashattention2) have reduced quadratic complexity to linear complexity at the cost of hiding attention scores, rendering many existing pruning methods obsolete. Addressing this challenge, SharpV pioneers a novel pruning paradigm that operates independently of exposed attention scores, providing fresh insights for visual token compression.
@@
As is well-known, cosine similarity effectively measures directional alignment between vectors. In this work, we aim to quantify token dissimilarity as a positive measure. While some approaches simply rescale cosine similarity using (1 - cosine similarity), this becomes problematic in high-dimensional spaces (cosdim1; cosdim2), where random vectors have expected cosine similarity approaching zero (cosine).
@@
While some approaches compute pairwise token similarity to preserve critical semantic information (divprune; llavascissor), this inevitably leads to quadratic computational complexity. Unlike existing methods (vicom2) that coarsely estimate frame-level uniqueness, we propose a novel token-level spatio-temporal importance assessment mechanism. Assisted by our designed Dissimilarity Computation Module, each token’s importance is comprehensively evaluated along dual dimensions.
@@
As shown in Fig. 3, we comprehensively compare the cosine similarity between hidden states $G_{l}=(S_{l},V_{l},I_{l})\in\mathbb{R}^{(N_{s}+N_{v}+N_{p})\times d}$ at different layers $l$ and their original positions in $G=(S,V,I)\in\mathbb{R}^{(N_{s}+N_{v}+N_{p})\times d}$ for both LLaVA-OneVision-7B and PLLaVA, visualizing their average values. A striking observation is that in both models, the visual tokens $V_{l}$ exhibit high similarity on average in shallow layers and demonstrate a clear decreasing trend to their original counterparts $V$ as layer depth increases. In contrast, system and instruction tokens rapidly converge to near-zero similarity (0-0.2 range). Mathematically, the similarity of random high-dimensional vectors approaches zero, suggesting that system and instruction tokens undergo more complex unknown transformations through Attention and MLP operations. Meanwhile, visual tokens show measurable degradation of original visual information via cosine similarity, which we term Visual Information Degradation. Although the exact relationships between transformed token embeddings remain unclear (simignore), the similarity patterns reveal that visual information undergoes rapid decay in shallow layers resembling human memory curves before stabilizing at low similarity levels (¡0.2) after layer 14. This suggests LLMs primarily process visual information in early layers, with diminishing importance in deeper layers.
@@
Research on visual attention reveals a consistent pattern: Inefficient Visual Attention phenomenon (fastv; vtw) validate that visual tokens receive drastically diminished attention scores in deeper layers of multimodal LLMs. We further quantify this via cosine similarity, a robust metric for high-dimensional feature decay (cosine2; cosine3; cosine4; cosine5).
From the perspective of information theory, this degradation pattern aligns with the concept of progressive information compression in hierarchical systems. The rapid similarity decay in early layers suggests intensive feature extraction, while the stable low similarity in deeper layers indicates visual information has been sufficiently encoded into subsequent representations. This mirrors the information bottleneck principle (bottleneck; videobottleneck), where networks discard irrelevant details while preserving task-relevant features through successive transformations (deepbottleneck). The distinct behavior of visual versus textual tokens further implies modality-specific information (crossmodal1; crossmodal2) processing pathways in multimodal LLMs, analogous to attentional load theory in human perception (visualattention1).
@@
We employ lmms-eval (lmms) as our primary evaluation framework. For PLLaVA (pllava) specifically, we use its official implementation to ensure fair comparison. Our comprehensive evaluation covers diverse video benchmarks, including MVBench (mvbench) and VideoMME (videomme) for multi-dimensional video understanding and NExT-QA (nextqa) and ActNet-QA (activitynet) for video question answering.
@@
We implement FastV (fastv), PruMerge (prumerge) and DyCoKe (dycoke) on two widely used LLMs for video understanding: PLLaVA(pllava) and LLaVA-OneVision (llavaov) with details shown in Tab.2. To demonstrate scalability, we evaluate 0.5B and 7B versions of LLaVA-OneVision.
@@
The term Token Budget combines the visual pruning ratio (VR) from the pre-LLM stage and the memory pruning ratio (MR) from the intra-LLM stage, defined as $VR\times MR$ following (dycoke; framefusion). VR measures the ratio of retained visual tokens fed into LLMs (vicom2; divprune), while MR estimates KV-Cache utilization during decoding (framefusion). For efficiency comparison, we evaluate FLOPs, GPU memory and acceleration metrics: Time To First Token (TTFT) and Time Per Output Token (TPOT).
@@ Conclusion
SharpV introduces a novel approach to intra-LLM visual token pruning that operates without reliance on exposed attention scores, achieving linear complexity while maintaining compatibility with Flash Attention. Furthermore, we propose a simple yet effective pre-LLM token selection method, offering a self-calibrated, adaptive thresholding scheme with strong generality and robustness based on information measurement. These components synergize to reduce the final pruning rate around 12%, while frequently improving accuracy across multiple benchmarks due to reduced visual noise, enhancing both efficiency and reliability in Video-LLM reasoning.